# Traversable Wormhole in AdS and Entanglement

### v1

@@ 1 Introduction
Like a black hole, a wormhole is a fascinating solution to Einstein’s equations that has garnered significant interest in the theoretical physics community. Notably, it plays a crucial role in the recent breakthroughs regarding the black hole information paradox [1, 2]. Generally, there are two types of wormholes: nontraversable and traversable. A well-known example of a nontraversable wormhole is the Einstein-Rosen bridge [3], which connects two asymptotic regions of eternal black holes. According to the “ER=EPR” conjecture [4], it is dual to entangled Einstein–Podolsky–Rosen (EPR) pairs, indicating a profound connection between wormholes and quantum entanglement. The traversable wormhole was first studied by Ellis [5] and Bronnikov [6] and then by Morris and Thorne [7]. However, these traversable wormholes generally violate the averaged null energy condition (ANEC), which raises questions about their existence. It’s important to note that the violation of ANEC can be accommodated through quantum fluctuations, such as the Casimir effect, which is considered acceptable. Nevertheless, the traversable wormhole described by Morris and Thorne allows for time travel to the past, which contradicts the law of causality [8]. The first traversable wormhole consistent with causality was proposed by Gao, Jafferis, and Wall using a double trace deformation [9]. In this model, it takes longer to travel through the wormhole throat than it does via the typical path outside. Thus, this traversable wormhole adheres to the principles of causality and may be realized in the physical world. A significant advancement in this area is the humanly traversable wormholes proposed by Maldacena, Milekhin, and Popov, which utilize Casimir-like energy from fermions or the Randall-Sundrum brane-world scenario [10, 11]. Following this, wormhole solutions to Einstein’s gravity coupled with Maxwell fields and two Dirac fermions were discovered [12] and refined in subsequent studies [13, 14]. Remarkably, this traversable wormhole is composed entirely of non-exotic matter.
@@ 3.1 Spherical topology
where a numerical cutoff of $x=\pm(1-\epsilon)$ should be taken. We take $\epsilon=10^{-4}$ in this section.
@@
we verify that the NEC is violated for our wormhole solution, i.e., $T_{\mu\nu}K^{\mu}K^{\nu}<0$. See Fig. 4. Thus, we indeed obtain a traversable wormhole in the asymptotically AdS spacetime.
@@ 3.2 Planar topology
We draw one typical solution of the planar wormhole and the charge density in Fig. 5. We also verify the violation of the NEC in Fig. 6, i.e., $T_{\mu\nu}K^{\mu}K^{\nu}<0$.
@@ 5.1 A toy model
Figure 12: The figure shows the disk radius has a minimal value (green point) for the connected phase. Besides, one disk radius corresponds to one disconnected surface and but two connected surfaces generally.

### v2

@@ 1 Introduction
Like a black hole, a wormhole is a fascinating solution to Einstein’s equations that has garnered significant interest in the theoretical physics community. Notably, it plays a crucial role in the recent breakthroughs regarding the black hole information paradox [1, 2]. Generally, there are two types of wormholes: nontraversable and traversable. A well-known example of a nontraversable wormhole is the Einstein-Rosen bridge [3], which connects two asymptotic regions of eternal black holes. According to the “ER=EPR” conjecture [4], it is dual to entangled Einstein–Podolsky–Rosen (EPR) pairs, indicating a profound connection between wormholes and quantum entanglement. The traversable wormhole was first studied by Ellis [5] and Bronnikov [6] and then by Morris and Thorne [7]. However, these traversable wormholes generally violate the averaged null energy condition (ANEC), which raises questions about their existence. It’s important to note that the violation of ANEC can be accommodated through quantum fluctuations, such as the Casimir effect, which is considered acceptable. Nevertheless, the traversable wormhole described by Morris and Thorne allows for time travel to the past, which contradicts the law of causality [8]. The first traversable wormhole consistent with causality was proposed by Gao, Jafferis, and Wall using a double trace deformation [9]. In this model, it takes longer to travel through the wormhole throat than it does via the typical path outside. Thus, this traversable wormhole adheres to the principles of causality and may be realized in the physical world. A significant advancement in this area is the humanly traversable wormholes proposed by Maldacena, Milekhin, and Popov, which utilize Casimir-like energy from fermions or the Randall-Sundrum brane-world scenario [10, 11]. Following this, wormhole solutions to Einstein’s gravity coupled with Maxwell fields and two Dirac fermions were discovered [12] and refined in subsequent studies [13, 14]. Remarkably, this traversable wormhole is composed entirely of non-exotic matter. See also [15, 16, 17, 18] for some recent developments of wormholes.
@@
Let us state the motivations for investigating the HEE in the wormhole background. First, HEE plays an essential role in AdS/CFT. It reveals a deep relation between the geometry and the quantum entanglement [20]. From the first law of entanglement entropy, one can derive the linear Einstein equations in the framework of AdS/CFT [21, 22]. Furthermore, the recent breakthrough in the black hole information problem also benefits from the developments of HEE, particularly the concept of entanglement islands [1, 2]. Second, “ER=EPR” shows that the nontraversable wormhole is closely related to the entanglement entropy [4]. Thus, it is interesting to study if there are also some relations between the traversable wormhole of this paper and the entanglement entropy. Third, the fermions on the two sides of the flat wormhole are entangled [12]. Thus, exploring the entanglement of two sides of the AdS wormhole is interesting, as we have done in this paper.
@@ 3.1 Spherical topology
where a numerical cutoff of $x=\pm(1-\epsilon)$ should be taken. We take $\epsilon=10^{-4}$ in this section. Note that we have $\lim_{x\rightarrow\pm 1}N(x)^{2}=\lim_{x\rightarrow\pm 1}B(x)^{2}\to\infty$ for the AdS wormholes (see Fig.1 ), while $\lim_{x\rightarrow\pm 1}N(x)^{2}=\lim_{x\rightarrow\pm 1}B(x)\to 1$ for the flat wormholes [12, 13, 14]. This is the main difference between our wormhole solutions and those presented in [12, 13, 14] for flat space.
@@
we verify that the NEC is violated for our wormhole solution, i.e., $T_{\mu\nu}K^{\mu}K^{\nu}<0$. See Fig. 4. Thus, we indeed obtain a traversable wormhole in the asymptotically AdS spacetime. Fig. 4 shows NEC is most severely violated near the throat of the wormhole $x\to 0$, while it is essentially not violated near the AdS boundary $x\to\pm 1$. More precisely, we numerically observe the NEC is violated in the bulk region $-0.67<x<1$.
@@ 3.2 Planar topology
We draw one typical solution of the planar wormhole and the charge density in Fig. 5. We also verify the violation of the NEC in Fig. 6, i.e., $T_{\mu\nu}K^{\mu}K^{\nu}<0$. We numerically find the NEC is violated in the bulk region $-0.76<x<1$ The closer to the wormhole throat, the more severely the NEC is violated.
@@ 4 HEE of strip
To end this section, let us comment on the possible relation between the NEC and the phase transition of holographic entanglement entropy. Recall that the NEC is most severely violated near the wormhole’s throat. Similarly, the HEE undergoes a phase transition when the RT surface tends deeply into the bulk region near the wormhole throat. Thus, it is natural to ask if violating the null energy condition causes the phase transition of holographic entanglement entropy. It is qualitatively correct. For instance, the turning point of the disconnected extremal surface at the phase-transition point is $\left|x_{\text{min}}\right|\approx 0.37$, localized in the region violating the NEC. However, we do not find an exact relation, if any, between the NEC and the phase transition of HEE. It is beyond the primary purpose of this paper. We leave this interesting open question for future work.
@@ 5.1 A toy model
Figure 12: The figure shows the disk radius has a minimal value (green point) for the connected phase. Besides, one disk radius corresponds to one disconnected surface and two connected surfaces generally.